OpenStudy (anonymous):

Find the value of Q in the following system so that the solution to the system is {x,y):x-3y=4} x-3y=4 Qx-6y=8

OpenStudy (anonymous):

i think i can help

OpenStudy (anonymous):

are you there?

OpenStudy (anonymous):

Seems that you may of left something out. Do you have the values of x and y that this particular system should give?

OpenStudy (anonymous):

No thats all that the problem gives /:

OpenStudy (anonymous):

Okay ill give it a go.

OpenStudy (anonymous):

Q = 2

OpenStudy (anonymous):

\[x-3y=4\] \[y = \frac{x-4}{3}\] \[Qx- 6(\frac{x-4}{3})= 8\] \[Qx -2(x-4) = 8\] \[Qx-2x+8=8\] \[x(Q-2) = 0\] \[Q-2=0\] \[Q=2\]

OpenStudy (anonymous):

Thank youu ! much appreciated(: